What are ΔE, q, w and ΔH for the evaporation of 10.3 g of Br2(l) at 298 K and 1 atm? (See Thermodynamic Properties.)
Determine ΔH, q, w, and ΔE at 298 K and 1 atm pressure for the reaction of 9.340 g of NH3(g) with excess HCl(g) to produce NH4Cl(s). (See Thermodynamic Properties.) Please use the mass of H at 1.0079 and the mass of N as 14.007. Use the values in the Thermodynamic Properties to determine the expected number of significant figures.

Please use the values in the Thermodynamic Properties to determine the expected number of significant figures.
Answer #1

Question 1.

What are ΔE, q, w and ΔH for the evaporation of 10.3 g of Br2(l) at 298 K and 1 atm? (See Thermodynamic Properties.)

dHvap = 30.91 kJ/mol

mol of Br2 = mass/MW = 10.3/159.808 = 0.06445 mol of Br2

H = n*dVap = 0.06445*30.91 = 1.99214 kJ

Q = H = 1.99214 kJ

P = 1 atm = 10325

Vf = form ideal gas, Vi = density

D = 3.1028 g/mL --> V = mass/D = 10.3/3.1028 = 3.319582 cm3 = 3.319582e-6*10^-6 m3

PV = nRT

V = nRT/P = (0.06445 *8.314)(298)/101325 = 0.0015759 m3

W = -P*(Vf-Vi) = -101325*(0.0015759- 3.319582*10^-6) = -159.34171J

dE = Q-W = 1.99214 *10^3 + 159.34171 = 2151.48171 J = 2.15 kJ